# Statistical Taylor Expansion111

### v7

@@ Abstract
Statistical Taylor expansion replaces the input precise variables in a conventional Taylor expansion with random variables each with known distribution, to calculate the result mean and deviation. It is based on the uncorrelated uncertainty assumption [1]: Each input variable is measured independently with fine enough statistical precision, so that their uncertainties are independent of each other. Unlike traditional statistics, it requires each input sample count as an input. Statistical Taylor expansion reveals that the intermediate analytic expressions can no longer be regarded as independent of each other, and the result of analytic expression should be path independent. This conclusion differs fundamentally from the conventional common approach in applied mathematics to find the best execution path for a result. This paper also presents an implementation of statistical Taylor expansion called variance arithmetic, and the tests on variance arithmetic.
@@ 2.6 Convergence
2.7 Bounding Factor
@@
figure The result uncertainty or precision (as shown by the y-axis on the left) for the different bounding factor $\kappa$ (as shown by the x-axis) for $(1\pm 0.05)^{-1}$, $(1\pm 0.1)^{-1}$, $(1\pm 0.1)^{2}$, $\log(1\pm 0.1)$, $e^{1\pm 0.1}$ and $\sin(0\pm 0.1)$. For better comparison, the precision of $(1\pm 0.05)^{-1}$ is normalized by $1/2$ while the precision of $(1\pm 0.1)^{2}$ is normalized by $2$. The bounding leakage (as shown by the y-axis on the right) for the different bounding factor $\kappa$ (as shown by the x-axis).
@@
Statistical Taylor expansion should be applied directly only when the bounding factor $\kappa$ is large enough. Figure 2.7 shows that due to bounding leakage, the result precision of $(1\pm 0.05)^{-1}$, $(1\pm 0.1)^{-1}$, $(1\pm 0.1)^{2}$, and $\log(1\pm 0.1)$, and the result uncertainty of $e^{1\pm 0.1}$ and $\sin(0\pm 0.1)$, are all less than the corresponding stable values when $4\lesssim\kappa$.
@@
In other cases, as a 1st-order approximation to guarantee $\zeta(2)$ to be $1$ for Formula (2.3), the result variance is normalized by $\zeta(2,\kappa)$. Figure 2.7 shows that the error after such adjustment is much less than those of Figure 2.7 (e.g., from about 70% to less than 0.5% for $x^{2}$ when $\kappa=1$), although the results of such adjustment vary for different functions (e.g., slightly smaller for $\sin(x)$ than $e^{x}$), different parameters (e.g., much less for $x^{2}$ than $x^{-1}$), and different input uncertainty (e.g., less for $(1\pm 0.05)^{-1}$ than $(1\pm 0.1)^{-1}$). Such adjustment is needed to obtain the approximate result for either $\log(x\pm\delta x)$ or $(x\pm\delta x)^{c}$ if the input precision is coarser than $1/4$, by deliberately selecting a small bounding range for statistical Taylor expansion.
2.8 Statistical Bounding
[…]
figure The measured bounding leakage (as shown by the y-axis) for the different bounding factor $\kappa$ (as shown by the x-axis) and sample count $N$ (as shown by the legend). The curve labeled sigma-rule corresponds to Formula (2.40) when $\varrho=\kappa$.
[…]
figure The measured bounding leakage precision (as shown by the y-axis) for the different bounding factor $\kappa$ (as shown by the x-axis) and sample count $N$ (as shown by the legend).
[…]
figure The measured bounding leakage (as shown by the y-axis on the left) and the bounding leakage precision (as shown by the y-axis on the right) for the different sample count $N$ (as shown by the x-axis). The bounding leakage precision decays to a constant about $0.95$ with $N^{-1.1}$.
[…]
figure The measured bounding factor $\kappa$ (as shown by the y-axis) for the different sample count $N$ (as shown by the x-axis) and the precision of $\epsilon(\kappa,N)$ (as shown by the legend). Also included is the relation $\zeta(2,\kappa)=N/(N-1)$, which is labeled ”Adjustment”.
With $\zeta(1)=0$, $\delta^{2}x=\zeta(2,\kappa)(\delta x)^{2}$ for $f(x)=x$. When $\kappa$ is bounded, $\zeta(2,\kappa)<1$, which leads to the following two possible forms of statistical Taylor expansion:
  • If $\zeta(n)=\zeta(n,\kappa)$ in Formula (2.3) and (2.3), $\delta^{2}x<(\delta x)^{2}$. As an explanatory model, $(\delta x)^{2}$ is the standard variance while $\delta^{2}x$ is the sample variance, so that $\zeta(2,\kappa)=(N-1)/N$ in which $N$ is the sample count. In this form, statistical Taylor expansion inputs standard variance, and outputs sample variance.
  • If $\zeta(n)=\zeta(n,\kappa)/\zeta(2,\kappa)$ in Formula (2.3) and (2.3), $\delta^{2}x=(\delta x)^{2}$. As shown in Figure 2.7, the result error of statistical Taylor expansion decreases exponentially with $\kappa$. In this form, statistical Taylor expansion inputs and outputs the same type of variance.
In addition to $\zeta(2,\kappa)=(N-1)/N$, the bounding leakage of sampling suggests another model of $\kappa(N)$.
[…]
$\simeq$  $\;1-\frac{1}{2}\xi(\frac{\varrho}{\sqrt{2}})-\frac{1}{2}\xi(\frac{\kappa}{\sqrt{2}});$ (2.40)
$\epsilon(\kappa,N)\equiv$  $\;1-P(-\varrho\leq\frac{x-\overline{x}}{\delta x}\leq\kappa);$ (2.41)
For an uncertainty distribution with density $\rho$, mean $\mu$, and deviation $\sigma$, Formula (2.39) shows the bounding leakage $\epsilon(\kappa)$. When the uncertainty distribution is Gaussian, Formula (2.39) becomes Formula (2.40), in which $\xi()$ is the Gaussian error function [5]. When $\varrho=\kappa$, Formula (2.40) is the familiar sigma-rule on $\kappa$ [5].
Similarly, Formula (2.41) shows the bounding leakage $\epsilon(\kappa,N)$ for sample mean $\overline{x}$ and sample deviation $\delta x$, which has the bounding range $[\overline{x}-\varrho\delta x,\overline{x}+\kappa\delta x]$. When sampling from normal distribution with $\varrho=\kappa$, the corresponding bounding leakage is $1-\frac{1}{2}\xi(\frac{1}{\sqrt{2}}|\kappa\delta x+\overline{x}|)-\frac{1}{2}\xi(\frac{1}{\sqrt{2}}|\kappa\delta x-\overline{x}|)$. Figure 2.8 shows that the measured bounding leakage $\lim_{N\rightarrow\infty}\epsilon(\kappa,N)=\epsilon(\kappa)$ with $\epsilon(\kappa)<\epsilon(\kappa,N)$. Figure 2.8 shows that the measured precision of $\epsilon(\kappa,N)$ decreases with $N$, and increases with $\kappa$ for each $N$. Figure 2.8 suggests that $N$ determines $\kappa$ and the result error in statistical Taylor expansion.
When the underlying uncertainty distribution is uniform, Figure 2.8 shows that while the bounding leakage $\epsilon(\kappa,N)$ decreases exponentially with the sample count $N$, the bounding leakage precision decreases to a constant near $0.95$ when $1000\leq N$. It suggests that the practical bounding leakage precision should be a constant close to $1$.
Figure 2.8 measured the relation between the sample count $N$ and the bounding factor $\kappa$ for the given bounding leakage precision of $0.368948671$, $0.5$, and $1$, in which $0.368948671$ is the measured bounding leakage precision for $N=2,\kappa=1$. When the bounding leakage precision is $1$, the bounding factor $\kappa$ decreases with $N$ when $N\leq 5$, which is undesirable. Such decreasing also exists when the bounding leakage precision is reduced to $0.5$, but to a less degree. Thus, $\kappa(N)$ in Figure 2.8 with the bounding leakage precision of $0.368948671$ could be another model for $\kappa(N)$, with $\kappa(N)\sim N^{0.229}$ when $10\ll N$. As a comparison, the model of $\zeta(2,\kappa)=(N-1)/N$ is labeled as ”Adjustment” in Figure 2.8, with $\kappa(N)\sim\log(N)$ when $10\ll N$. Which trend is preferred is a new question for statistics.
@@ 2.9 Dependency Tracing
    $\zeta(2n+1)=0:$
    $\;\;\;\delta^{2}\left(\frac{f^{(m)}_{x}\tilde{x}^{m}}{m!}+\frac{f^{(n)}_{x}\tilde{x}^{n}}{n!}\right)=(\delta x)^{2m}(\frac{f^{(m)}_{x}}{m!})^{2}\eta(2m)+(\delta x)^{2n}(\frac{f^{(n)}_{x}}{n!})^{2}\eta(2n)$
@@ 10.3 Improvement Needed
  • How to extend variance arithmetic to other input uncertainty distribution? Particular, when the underlying distribution is uniform, the analysis based on $\kappa$ no longer holds.

### v8

@@ Abstract
Statistical Taylor expansion replaces the input precise variables in a conventional Taylor expansion with random variables each with known distribution, to calculate the result mean and deviation. It is based on the uncorrelated uncertainty assumption [1]: Each input variable is measured independently with fine enough statistical precision, so that their uncertainties are independent of each other. It requires each sampling count to be sufficiently large for accurate results. Statistical Taylor expansion reveals that the intermediate analytic expressions can no longer be regarded as independent of each other, and the result of analytic expression should be path independent. This conclusion differs fundamentally from the conventional common approach in applied mathematics to find the best execution path for a result. This paper also presents an implementation of statistical Taylor expansion called variance arithmetic, and the tests on variance arithmetic.
@@ 2.6 Convergence
2.7 Bounding Factor and Sample Count
@@
figure The measured bounding leakage (as shown by the y-axis) for the different measuring bounding factor $\kappa_{s}$ (as shown by the x-axis) and sample count $N$ (as shown by the legend).
[…]
figure The measured bounding factor $\kappa_{r}$ (as shown by the y-axis on the left) and the corresponding measured bounding leakage $\epsilon(\kappa_{r})$ (as shown by the y-axis on the right) for the different sample count $N$ (as shown by the x-axis) when the underlying distribution is either uniform or normal (as shown by the legend, with different measuring bounding factor for the normal distribution).
$\epsilon(\kappa)=$  $\;1-\frac{1}{2}\xi(\frac{\varrho}{\sqrt{2}})-\frac{1}{2}\xi(\frac{\kappa}{\sqrt{2}});$ (2.39)
$\epsilon(\kappa,N)=$  $\;1-\frac{1}{2}\xi(\frac{|\varrho\delta x-\overline{x}|}{\sqrt{2}})-\frac{1}{2}\xi(\frac{|\kappa\delta x+\overline{x}|}{\sqrt{2}});$ (2.40)
When sampling from a underlying distribution with the mean $\mu$ and deviation $\sigma$, the sample mean $\overline{x}$ and sample deviation $\delta x$ approaches $\mu$ and $\sigma$ respectively with the sample count $N$ as $\sqrt{N}$, so that the sample range $[\overline{x}-\kappa\delta x,\overline{x}+\kappa\delta x]$ is different from the actual range $[\mu-\kappa\sigma,\mu+\kappa\sigma]$, to have a different bounding leakage $\epsilon(\kappa,N)$ vs. $\epsilon(\kappa)$.
[…]
Generally, $\kappa$ in Formula (2.4) increases with the sample count $N$, which has to be sufficiently large to get accurate result using statistical Taylor expansion.
2.8 Approximation for Insufficient Sample Count
[…]
figure The result uncertainty or precision (as shown by the y-axis on the left) and the bounding leakage (as shown by the y-axis on the right) for the different bounding factor $\kappa$ (as shown by the x-axis) for $(1\pm 0.05)^{-1}$, $(1\pm 0.1)^{-1}$, $(1\pm 0.1)^{2}$, $\log(1\pm 0.1)$, $e^{1\pm 0.1}$ and $\sin(0\pm 0.1)$. For better comparison, the precision of $(1\pm 0.05)^{-1}$ is normalized by $1/2$ while the precision of $(1\pm 0.1)^{2}$ is normalized by $2$.
@@ 2.8 Approximation for Insufficient Sample Count
When $N$ or $\kappa$ is not sufficiently large, the result variance of statistical Taylor may be too small due to too much bounding leakage. Figure 2.8 shows that the result precision of $(1\pm 0.05)^{-1}$, $(1\pm 0.1)^{-1}$, $(1\pm 0.1)^{2}$, and $\log(1\pm 0.1)$, and the result uncertainty of $e^{1\pm 0.1}$ and $\sin(0\pm 0.1)$, are all less than the corresponding stable values when $\kappa=5$ for an input with Gaussian uncertainty. To compensate for this effect, the result can be normalized by $\zeta(2,\kappa)$. Figure 2.8 shows that the error after such normalization is much less than those of Figure 2.8 (e.g., from about 70% to less than 0.5% for $x^{2}$ when $\kappa=1$), although the results of such adjustment vary for different functions (e.g., smaller for $\sin(x)$ than $\log(x)$), different parameters (e.g., much less for $x^{2}$ than $x^{-1}$), and different input uncertainty (e.g., less for $(1\pm 0.05)^{-1}$ than $(1\pm 0.1)^{-1}$), so the result of such normalization needs to be analyzed case by case.
On the other hand, either $\log(x\pm\delta x)$ or $(x\pm\delta x)^{c}$ will not converge if the input $P(x)$ is larger than $1/\kappa$. To obtain the result, an artificial $\kappa<1/P(x)$ can be used then normalized to obtain an approximate result.
@@ 2.9 Dependency Tracing
$\delta^{2}\left(\frac{f^{(m)}_{x}\tilde{x}^{m}}{m!}+\frac{f^{(n)}_{x}\tilde{x}^{n}}{n!}\right)=$ $\;(\delta x)^{2m}(\frac{f^{(m)}_{x}}{m!})^{2}\eta(2m)+(\delta x)^{2n}(\frac{f^{(n)}_{x}}{n!})^{2}\eta(2n)$ (2.45)
@@ 10.3 Improvement Needed
  • Why there is the discrepancy of measured $\kappa$ vs calculated $\kappa$ in Figure LABEL:fig:_Uniform_Bounding_Factor?
  • When the sample counts are not sufficient, how to normalize the result variance of the statistical Taylor expansion of $f(x,y,...)$ as in Figure LABEL:fig:_Normal_Bounding_Factor?